# Stepwise Regression and Pre-trained Edge for Robust Stereo Matching

### v2

@@ Abstract
Due to the difficulty in obtaining real samples and ground truth, the generalization performance and the fine-tuned performance are critical for the feasibility of stereo matching methods in real-world applications. However, the presence of substantial disparity distributions and density variations across different datasets presents significant challenges for the generalization and fine-tuning of the model. In this paper, we propose a novel stereo matching method, called SR-Stereo, which mitigates the distributional differences across different datasets by predicting the disparity clips and uses a loss weight related to the regression target scale to improve the accuracy of the disparity clips. Moreover, this stepwise regression architecture can be easily extended to existing iteration-based methods to improve the performance without changing the structure. In addition, to mitigate the edge blurring of the fine-tuned model on sparse ground truth, we propose Domain Adaptation Based on Pre-trained Edges (DAPE). Specifically, we use the predicted disparity and RGB image to estimate the edge map of the target domain image. The edge map is filtered to generate edge map background pseudo-labels, which together with the sparse ground truth disparity on the target domain are used as a supervision to jointly fine-tune the pre-trained stereo matching model. These proposed methods are extensively evaluated on SceneFlow, KITTI, Middbury 2014 and ETH3D. The SR-Stereo achieves competitive disparity estimation performance and state-of-the-art cross-domain generalisation performance. Meanwhile, the proposed DAPE significantly improves the disparity estimation performance of fine-tuned models, especially in the textureless and detail regions.
@@ I Introduction
Depth information is key to computer vision and graphics research in the real world. Accurate depth estimation is vital for fields such as autonomous driving, robot navigation and action recognition. Common depth estimation methods include structured light [1], stereo matching [2, 3, 4], radar [5], etc. The structured light method projects a coded pattern onto the object surface and estimates the depth by observing the distortion of the pattern imaged on the object surface. Due to the need to project multiple patterns consecutively, structured light methods are generally only suitable for static indoor scenes. The radar obtains a sparse depth map by transmitting pulses and receiving echoes. Millimeter-wave radar has a long detection range but fuzzy spatial resolution, while laser radar has a high spatial resolution but is expensive. In contrast, stereo matching calculates the depth map by estimating the horizontal displacement map (i.e., the disparity) of the pixels between the corrected left and right image pairs, and is therefore considered a dense depth estimation method with wide applicability and low cost.
@@
Many learning-based stereo methods [12, 6, 7, 8, 9, 11, 10] have achieved encouraging success. Some works [12, 13, 14] focus on designing the 4D cost volume, which characterizes the cost of pixel matching between left and right image pairs. They use 3D convolutional networks to aggregate and regularize the entire cost volume, and regress it to obtain the disparity. However, the large number of 3D convolutions leads to large computational and memory cost, which limits the applicability of such methods. Recently, iteration-based methods [16, 15, 17, 22, 20, 18, 21, 19] have shown great potential in terms of both accuracy and real-time performance, and thus have become the mainstay of current research. Iterative-based methods index the similarity features within current disparity neighborhood to the cost volume and then use it to update the disparity. This approach avoids the expensive cost of aggregating and regularizing cost volume and achieve a balance between performance and efficiency by controlling the number of iterations. In addition, some state-of-the-art iteration-based methods [17, 18] utilize a lightweight 3D convolutional network to regress the coarse initial disparity, thus further accelerating the iteration efficiency.
@@
However, stereo methods usually suffer from the problems of few real samples and difficult access to ground truth in practical applications. Existing methods [26, 25, 27, 24, 23] are generally pre-trained on a virtual dataset (e.g., SceneFlow [29]) and fine-tuned or directly inferred on real datasets, which imposes higher requirement on the generalization performance. In addition, there are significant differences in disparity distribution and density among different datasets due to variations in acquisition equipment and environment. The Middlebury 2014 [30], which uses hand labeling to obtain dense ground truth, has a maximum disparity value of more than 250 pixels. The KITTI dataset  [32, 31], which uses radar and careful calibration to obtain sparse ground truth, has all of its disparities within 100 pixels. These differences make it difficult for a method that performs well on one dataset to directly achieve the same performance on other datasets. Therefore, we need to improve the method so that it has good performance in estimating disparity across domains.
@@
In this paper, we claim that by designing a regression target with a similar distribution across different datasets, excellent performance can be consistently achieved using the same method across various datasets. Therefore, a disparity stepwise regression (DSR) is proposed to study the disparity estimation in robust stereo matching. Specifically, the regression target in the DSR is multiple disparity clips, which have the same range across different datasets. By summing the multiple disparity clips, the DSR enables accurate and robust disparity estimation. Meanwhile, a loss weight related to the regression target scale is proposed to improve the accuracy of disparity clips because it effectively mitigates the long-tailed distribution problem [33, 34] of the disparity clips. Based on these two simple but powerful methods, we design a novel architecture, Stepwise Regression Stereo (SR-Stereo), for robust disparity estimation.
@@
Furthermore, we observe that models fine-tuned with sparse ground truth suffer from severe edge blurring, i.e., edge disparity errors. As shown in Fig. 1, even the state-of-the-art models [16, 17] in the KITTI online leaderboard fail to accurately estimate the edge disparity. This further limits the effectiveness of stereo methods in real-world applications, as the vast majority of efficient labeling methods (e.g., radar) can only obtain sparse ground truth. The most direct way to mitigate this problem is to introduce edge-dependent supervision, but current edge feature extraction methods [35, 36, 37, 38, 39, 40] have difficulty in obtaining accurate and concise edges as pseudo-label due to the overly sparse disparity ground truth and the interference of object textures in RGB images.
@@
In this paper, a Domain Adaptation Based on Pre-trained Edge (DAPE) is proposed to address the issue of blurring in edge details. The two key observations behind this approach are: 1) disparity has more concise and reliable information about object edges compared to RGB maps. 2) models pre-trained with dense ground truth tend to have more accurate edge in predicted disparity on new domains. Therefore, we first propose a lightweight edge estimator that takes both disparity and RGB image as inputs to generate a dense edge map. This estimator, along with SR-Stereo, is pre-trained on a virtual dataset. Then, we utilize the pre-trained SR-Stereo to generate the disparity map for the target domain, which is further fed into the edge estimator to predict the edge map. In this process, the background pixels are retained and utilized as edge pseudo-label, while the foreground pixels are filtered out. Finally, we use the edge pseudo-label as additional supervision during the model fine-tuning process, which effectively improves the model’s disparity estimation performance on edge details.
@@
  • We propose a novel architecture, called SR-Stereo, which predicts disparity clips to alleviate distribution discrepancies across different datasets and uses a loss weight related to the regression target scale to improve the accuracy of the disparity clips. Experiments on RAFT-Stereo and IGEV-Stereo also demonstrate that the proposed stepwise regression architecture has better generalization performance than iteration-based methods.
  • We perform experiments on SceneFlow, KITTI, Middlebury 2014 and ETH3D. Compared to existing methods, the proposed SR-Stereo achieves the best cross-domain generalization performance across multiple datasets and obtains competitive results on the SceneFlow benchmark.
@@ II-A Iterative-based Stereo Matching
Compared to cost aggregation-based methods [41, 42, 43], iteration-based methods achieve significant improvement in both efficiency and accuracy. RAFT-Stereo [44, 16] is the first iteration-based method that constructs cost volume by calculating the dot product of features with same height. At the same time, it introduces multi-level GRU unit [45] which updates the disparity by indexing the features of cost volume. Based on RAFT-Stereo, DLNR [22] uses decoupled LSTM [46] instead of GRU to retain more high-frequency information during the iteration process, and designs a normalized refinement module to capture more detailed information at full resolution. The refinement of high-frequency information effectively improves the performance of disparity estimation. Further, CREStereo [15] applies the iterative process to different resolutions of disparity. It uses a deformable search window to more accurately index the features of cost volume, and proposes a hierarchical refinement network to update the disparity at different resolutions in a coarse-to-fine order. These methods focus more on improving the iterative process. In contrast, IGEV-Stereo [17] constructs a geometric encoding volume by combining Group-wise correlation and All-pairs Correlation, and uses a lightweight 3D regularization network to regress the initial disparity, which effectively improves the iterative efficiency and accuracy.
Our proposed SR-Stereo is similar in process to the iteration-based methods, but has a different target. The iteration-based methods focus on how to efficiently converge the disparity, and thus all of its update units target the ground truth disparity. Instead, the proposed SR-Stereo focuses on improving the generalization performance by mitigating the difference in the disparity distribution among different domains. To this end, the update units of SR-Stereo are all individually assigned different targets, i.e., range-constrained disparity clips.
@@ II-B Robust Stereo Matching
Since real samples and ground truth disparity are difficult to obtain for practical applications, some works [26, 27, 28] focus on cross-domain generalization, which aims to ensure good performance of the model on unseen datasets. CFNet [26] and UCFNet [27] narrow the domain difference by cascaded cost volume and multi-stage refinement of disparity. They also adaptively adjust the search space of disparity at each stage by uncertainty estimation. DSMNet [28] employs domain normalization and a trainable structure-preserving graph-based filter to extract robust geometric feature to improve the cross-domain generalization performance. These methods enhance the generalization performance by improving the network structure. In contrast, our proposed DSR designs a new regression target with similar distribution across different datasets to reduce the generalization difficulty.
@@ III-A Framework Overview
In this paper, we propose a novel stereo matching architecture, SR-Stereo, which achieves robust disparity estimation by stepwise regression of the disparity. In addition, we provide a general and reliable framework, DAPE, for model fine-tuning on sparse ground truth by digging into the disparity-based edge estimation. As shown in Fig. 2, the proposed DAPE consists of three steps:
@@
    Firstly, a robust stereo model and a lightweight edge estimator are pre-trained on a large synthetic dataset with dense ground truth. Specifically, a stepwise regression stereo (SR-Stereo) is proposed to mitigate the disparity distribution difference between different datasets. Meanwhile, the disparities predicted by SR-Stereo are fed jointly with the RGB images into a lightweight network to learn the edge information.
@@ III-B Stepwise Regression Stereo
Figure 3: The overall architecture of the proposed SR-Stereo. Compared to iteration-based methods, SR-Stereo is specially designed in terms of the update unit and the regression target. Specifically, a stepwise regression unit is proposed to control the range of disparity clips. Accordingly, a regression target segmentation is proposed to design a separate regression target for each stepwise regression unit.
Fig. 3 shows the overall architecture of the proposed SR-Stereo, which is improved based on the iterative stereo method IGEV-Stereo. Intuitively, the proposed SR-Stereo reconstructs the update unit and the regression target, and proposes a loss weight related to the regression target scale to improve the accuracy of disparity clips.
@@ III-B1 Feature Encoder
Following IGEV-Stereo, we use MobileNetV2 [47] pre-trained on ImageNet [48] and a series of upsampling blocks to extract multi-scale features $f_{1}^{l(r),i}\in\mathbb{R}^{C_{i}\times H/2^{i+1}\times W/2^{i+1}}$ ($i$=1,2,3,4 and $C_{i}$=48,64,192,160) of the left and right images $I_{l(r)}\in\mathbb{R}^{3\times H\times W}$.
@@ III-B2 Context Encoder
We use a series of residual blocks [49] and downsampling layers to extract multi-scale context features $f_{2}^{j}\in\mathbb{R}^{C_{j}\times H/2^{j+1}\times W/2^{j+1}}$ ($j$=1,2,3,4 and $C_{j}$=128) from the left image. These features are inserted into the stepwise regression unit to provide global information.
@@ III-B3 Fused Cost Volume
By combining Group-wise correlation and Correlation, we construct the cost volume $G_{c}$ based on $f_{1}^{l(r),1}$. Then, we use a lightweight 3D UNet to aggregate the cost volume and obtain the initial disparity $d_{init}\in\mathbb{R}^{1\times H/4\times W/4}$.
@@ III-B4 Disparity Stepwise Regression
The variation in disparity distribution across different domains poses a challenge to the generalization performance of existing stereo matching methods. In order to address this issue, we propose a disparity stepwise regression (DSR), which obtains the disparity by predicting multiple disparity clips with the same range. It consists of two parts, including stepwise regression unit and regression target segmentation.
@@
Stepwise Regression Unit. Fig. 4 shows the architecture of the stepwise regression unit. For each regression, we index a set of features $f_{G}$ from the cost volume $G_{c}$ using the cumulative disparity-centered set as follows:
@@
where $d_{k}$ is the cumulative disparity ($d_{0}$ = $d_{init}$) and $p$ denotes the average pooling operation. These features reflect the matching cost within the neighborhood of $d_{k}$. We use them along with the cumulative disparity $d_{k}$ to update the hidden state $h_{k}$ of ConvGRU [17]:
$h_{k+1}=ConvGRU(f_{G},d_{k},h_{k},f_{2})$ (2)
@@
$d_{k+1}^{full}=Upsampling(f_{1}^{l,i},I_{l},4d_{k+1})$ (6)
Regression Target Segmentation. We specify how to obtain the ground truth disparity clip $\Delta d_{gt,k}$ and the full-resolution ground truth cumulative disparity $d_{gt,k+1}^{full}$. Specifically, the ground truth of the full-resolution cumulative disparity is constantly changing with the stepwise regression process, as shown below:
@@
where $clip$ denotes the clipping operation and $d_{gt}$ is the ground truth disparity. In addition, we use linear interpolation to generate the ground truth disparity clip $\Delta d_{gt,k}$:

### v3

@@ Abstract
Due to the difficulty in obtaining real samples and ground truth, the generalization performance and the fine-tuned performance are critical for the feasibility of stereo matching methods in real-world applications. However, the diverse datasets exhibit substantial discrepancies in disparity distribution and density, thus presenting a formidable challenge to the generalization and fine-tuning of the model. In this paper, we propose a novel stereo matching method, called SR-Stereo, which mitigates the distributional discrepancies across different datasets by predicting the disparity clips and uses a loss weight related to the regression objective scale to improve the accuracy of the disparity clips. Moreover, this stepwise regression architecture can be easily extended to existing iteration-based methods to improve the performance without changing the structure. In addition, to mitigate the edge blurring of the fine-tuned model on sparse ground truth, we propose Domain Adaptation Based on Pre-trained Edge (DAPE). Specifically, we use the predicted disparity and RGB image to estimate the edge map of the target domain image. The edge map is filtered to generate edge map background pseudo-labels, which together with the sparse ground truth disparity on the target domain are used as a supervision to jointly fine-tune the pre-trained stereo matching model. These proposed methods are extensively evaluated on SceneFlow, KITTI, Middbury 2014 and ETH3D. The SR-Stereo achieves competitive disparity estimation performance and state-of-the-art cross-domain generalisation performance. Meanwhile, the proposed DAPE significantly improves the disparity estimation performance of fine-tuned models, especially in the textureless and detail regions.
@@ I Introduction
Depth information is key to computer vision and graphics research in the real world. Accurate depth estimation is vital for fields such as autonomous driving, robot navigation and action recognition. Common depth estimation methods include structured light [1], stereo matching [2, 3, 4], radar [5], etc. The structured light method projects a coded pattern onto the object surface and estimates the depth by observing the distortion of the pattern imaged on the object surface. Due to the need to project multiple patterns consecutively, structured light methods are generally only suitable for static indoor scenes. The radar obtains a sparse depth map by transmitting pulses and receiving echoes. Millimeter-wave radar has a long detection range but fuzzy spatial resolution, while laser radar has a high spatial resolution but is expensive. In contrast, stereo matching calculates the depth map by estimating the horizontal displacement map (i.e., the disparity) of the pixels between the corrected left and right image pairs. By utilizing only two cameras, stereo matching is able to provide dense depth maps in different scenes, and is therefore considered a cost-effective and widely applicable method for depth estimation.
@@
Many learning-based stereo methods [12, 6, 7, 8, 9, 11, 10] have achieved encouraging success. Some works [12, 13, 14] focus on designing the 4D cost volume, which characterizes the cost of pixel matching between left and right image pairs. They use 3D convolutional networks to aggregate and regularize the entire cost volume, and regress it to obtain the disparity. However, the large number of 3D convolutions leads to large computational and memory cost, which limits the applicability of such methods. Recently, iteration-based methods [16, 15, 17, 22, 20, 18, 21, 19] have shown great potential in terms of both accuracy and real-time performance, and thus have become the mainstay of current research. Iterative-based methods index the similarity features within current disparity neighborhood from the cost volume and then use it to update the disparity. This approach avoids the expensive cost of aggregating and regularizing cost volume and achieves a balance between performance and efficiency by controlling the number of iterations. In addition, some state-of-the-art iteration-based methods [17, 18] utilize a lightweight 3D convolutional network to regress the coarse initial disparity, thus further accelerating the iteration efficiency.
@@
However, stereo methods usually suffer from the problems of few real samples and difficult access to ground truth in practical applications. Existing methods [26, 25, 27, 24, 23] are generally pre-trained on a virtual dataset (e.g., SceneFlow [29]) and fine-tuned or directly inferred on real datasets, which imposes higher requirement on the generalization performance. In addition, there are significant discrepancies in disparity distribution and density among different datasets due to variations in acquisition equipment and environment. The Middlebury 2014 [30], which uses hand labeling to obtain dense ground truth, has a maximum disparity value of more than 250 pixels. The KITTI dataset  [32, 31], which uses radar and careful calibration to obtain sparse ground truth, has all of its disparities within 100 pixels. These discrepancies make it difficult for a method that performs well on one dataset to directly achieve the same performance on other datasets. Therefore, we need to improve the method so that it has good performance in estimating disparity across domains.
@@
In this paper, we claim that by designing a regression objective with a similar distribution across different datasets, excellent performance can be consistently achieved using the same method across various datasets. Therefore, a disparity stepwise regression (DSR) is proposed to study the disparity estimation in robust stereo matching. Specifically, the regression objective in the DSR is multiple disparity clips, which have the same range across different datasets. By summing the multiple disparity clips, the DSR enables accurate and robust disparity estimation. Meanwhile, a loss weight related to the regression objective scale is proposed to improve the accuracy of disparity clips because it effectively mitigates the long-tailed distribution problem [33, 34] of the disparity clips. Based on these two simple but powerful methods, we design a novel architecture, Stepwise Regression Stereo (SR-Stereo), for robust disparity estimation.
@@
Furthermore, we observe that models fine-tuned with sparse ground truth suffer from severe edge blurring, i.e., edge disparity errors. As shown in Figure 1, even the state-of-the-art models [16, 17] in the KITTI online leaderboard fail to accurately estimate the edge disparity. This further limits the effectiveness of stereo methods in real-world applications, as the vast majority of efficient labeling methods (e.g., radar) can only obtain sparse ground truth. The most direct way to mitigate this problem is to introduce edge-dependent supervision, but current edge feature extraction methods [35, 36, 37, 38, 39, 40] have difficulty in obtaining accurate and concise edges as pseudo-label due to the overly sparse disparity ground truth and the interference of object textures in RGB images.
@@
In this paper, a Domain Adaptation Based on Pre-trained Edge (DAPE) is proposed to address the issue of blurring in edge details. The two key observations behind this approach are: 1) disparity has more concise and reliable information about object edges compared to RGB maps. 2) models pre-trained with dense ground truth tend to have accurate edge in predicted disparity on new domains. Therefore, we first propose a lightweight edge estimator that takes both disparity and RGB image as inputs to generate a dense edge map. This estimator, along with SR-Stereo, is pre-trained on a virtual dataset. We utilize the pre-trained SR-Stereo to generate the disparity map for the target domain, which is further fed into the edge estimator to predict the edge map. Then, the background pixels in the edge map are retained and utilized as edge pseudo-label, while the foreground pixels are filtered out. Finally, we use the edge pseudo-label as additional supervision during the model fine-tuning process, which effectively improves the model’s disparity estimation performance on edge details.
@@
  • We propose a novel architecture, called SR-Stereo, which predicts disparity clips to alleviate distribution discrepancies across different datasets and uses a loss weight related to the regression objective scale to improve the accuracy of the disparity clips. Experiments on RAFT-Stereo and IGEV-Stereo also demonstrate that the proposed stepwise regression architecture has better generalization performance than iteration-based methods.
@@
  • We perform experiments on SceneFlow, KITTI, Middlebury 2014 and ETH3D. Compared to existing methods, the proposed SR-Stereo achieves the state-of-the-art cross-domain generalization performance across multiple datasets and obtains competitive results on the SceneFlow and KITTI benchmark. Meanwhile, experimental results on multiple realistic datasets show that the proposed DAPE significantly improves the disparity estimation performance of the fine-tuned model, especially in untextured and detailed regions.
@@ II-A Iterative-based Stereo Matching
Compared to cost aggregation-based methods [41, 42, 43], iteration-based methods achieve significant improvement in both efficiency and accuracy. RAFT-Stereo [44, 16] is the first iteration-based method that constructs cost volume by calculating the dot product of features with same height. At the same time, it introduces multi-level GRU (Gated Recurrent Unit) unit [45] which updates the disparity by indexing the features of cost volume. Based on RAFT-Stereo, DLNR [22] uses decoupled LSTM [46] (Long Short-Term Memory) instead of GRU to retain more high-frequency information during the iteration process, and designs a normalized refinement module to capture more detailed information at full resolution. The refinement of high-frequency information effectively improves the performance of disparity estimation. Further, CREStereo [15] applies the iterative process to different resolutions of disparity. It uses a deformable search window to more accurately index the features of cost volume, and proposes a hierarchical refinement network to update the disparity at different resolutions in a coarse-to-fine order. In recent research, cost aggregation based on 3D convolutional networks is revisited. For instance, IGEV-Stereo [17] constructs a geometric encoding volume by combining Group-wise correlation and All-pairs Correlation, and uses a lightweight 3D regularization network to regress the initial disparity, which effectively improves the iterative efficiency and accuracy.
Our proposed SR-Stereo is similar in process to the iteration-based methods, but has a different objective. The iteration-based methods focus on how to efficiently converge the disparity, and thus all of its update units target the ground truth disparity. Instead, the proposed SR-Stereo focuses on improving the generalization performance by mitigating the discrepancies in the disparity distribution among different domains. To this end, the update units of SR-Stereo are all individually assigned different objectives, i.e., range-constrained disparity clips.
@@ II-B Robust Stereo Matching
Since real samples and ground truth disparity are difficult to obtain for practical applications, some works [26, 27, 28] focus on cross-domain generalization, which aims to ensure good performance of the model on unseen datasets. CFNet [26] and UCFNet [27] narrow the domain discrepancy by cascaded cost volume and multi-stage refinement of disparity. They also adaptively adjust the search space of disparity at each stage by uncertainty estimation. DSMNet [28] employs domain normalization and a trainable structure-preserving graph-based filter to extract robust geometric feature to improve the cross-domain generalization performance. These methods enhance the generalization performance by improving the network structure. In contrast, our proposed DSR designs a new regression objective with similar distribution across different datasets to reduce the generalization difficulty.
@@ III-A Framework Overview
In this paper, we propose a novel stereo matching architecture, SR-Stereo, which achieves robust disparity estimation by stepwise regression of the disparity. In addition, we provide a general and reliable framework, DAPE, for model fine-tuning on sparse ground truth by digging into the disparity-based edge estimation. As shown in Figure 2, the proposed DAPE consists of three steps:
@@
    Firstly, a robust stereo model and a lightweight edge estimator are pre-trained on a large synthetic dataset with dense ground truth. Specifically, a stepwise regression stereo (SR-Stereo) is proposed to mitigate the disparity distribution discrepancies between different datasets. Meanwhile, the disparities predicted by SR-Stereo are fed jointly with the RGB images into a lightweight network to learn the edge information.
@@ III-B Stepwise Regression Stereo
Figure 3: The overall architecture of the proposed SR-Stereo. Compared to iteration-based methods, SR-Stereo is specially designed in terms of the update unit and the regression objective. Specifically, a stepwise regression unit is proposed to control the range of disparity clips. Accordingly, a regression objective segmentation is proposed to design a separate regression objective for each stepwise regression unit.
Figure 3 shows the overall architecture of the proposed SR-Stereo, which is improved based on the iterative stereo method IGEV-Stereo. Intuitively, the proposed SR-Stereo reconstructs the update unit and the regression objective, and proposes a loss weight related to the regression objective scale to improve the accuracy of disparity clips.
@@ III-B1 Feature Encoder
Following IGEV-Stereo, we use MobileNetV2 [47] pre-trained on ImageNet [48] and a series of upsampling blocks to extract multi-scale features $f_{1}^{left,i}\in\mathbb{R}^{C_{i}\times H/2^{i+1}\times W/2^{i+1}}$ from the left image $I_{left}\in\mathbb{R}^{3\times H\times W}$ and $f_{1}^{right,i}\in\mathbb{R}^{C_{i}\times H/2^{i+1}\times W/2^{i+1}}$ from the right image $I_{right}\in\mathbb{R}^{3\times H\times W}$, respectively. The $i$=1,2,3,4 and the feature size $C_{i}$=48,64,192,160.
@@ III-B2 Context Encoder
We use a series of residual blocks [49] and downsampling layers to extract multi-scale context features $f_{2}^{left,j}\in\mathbb{R}^{C_{j}\times H/2^{j+1}\times W/2^{j+1}}$ ($j$=1,2,3,4 and $C_{j}$=128) from the left image. These features are inserted into the stepwise regression unit to provide global information.
@@ III-B3 Fused Cost Volume
By combining Group-wise correlation and Correlation, we construct the cost volume $G_{c}$ based on $f_{1}^{left,1}$ and $f_{1}^{right,1}$. Then, we use a lightweight 3D UNet to aggregate the cost volume and obtain the initial disparity $d_{init}\in\mathbb{R}^{1\times H/4\times W/4}$.
@@ III-B4 Disparity Stepwise Regression
We propose a disparity stepwise regression, which obtains the disparity by predicting multiple disparity clips with the same range. It consists of two parts, including stepwise regression unit and regression objective segmentation.
@@
Stepwise Regression Unit. Figure 4 shows the architecture of the stepwise regression unit. For each regression, we index a set of features $f_{G}$ from the cost volume $G_{c}$ using the cumulative disparity-centered set as follows:
@@
where $d_{k}$ is the cumulative disparity ($d_{0}$ = $d_{init}$) and $p$ denotes the average pooling operation. We use these features from $G_{c}$ along with the cumulative disparity $d_{k}$ to update the hidden state $h_{k}$ of ConvGRU [17]:
$h_{k+1}=ConvGRU(f_{G},d_{k},h_{k},f_{2}^{left})$ (2)
@@
$d_{k+1}^{full}=Upsampling(f_{1}^{left,i},I_{left},4d_{k+1})$ (6)
Regression Objective Segmentation. We specify how to obtain the ground truth disparity clip $\Delta d_{gt,k}$ and the full-resolution ground truth cumulative disparity $d_{gt,k+1}^{full}$. Specifically, the ground truth of the full-resolution cumulative disparity is constantly changing with the stepwise regression process, as shown below:
@@
in which
$clip_{-M}^{M}(x)=\begin{cases}-M&\text{ if }x<-M\\ x&\text{ if }-M\leq x\leq M\\ M&\text{ if }x>M\end{cases}$ (8)
where $d_{gt}$ is the ground truth disparity and $clip$ denotes the clipping operation.
In addition, we use linear interpolation to generate the ground truth disparity clip $\Delta d_{gt,k}$: